### v2

@@ 4.4 Competition between Searchers
In the previous section, we highlighted the importance of competition between Rsync-bot and Titan-bot. It turns out that these two searchers compete intensely throughout our entire study period. And, like in auction for block 21,322,649, their competition is a key determinant of which builder wins the block: every auction cycle for which we have data includes at least one transaction from these bots, which together account for approximately 60% of the builders’ revenues in the blocks where they appear, underscoring their central role in shaping auction outcomes.
@@
We identify 57 instances in which Titan-bot and Rsync-bot compete for execution in the same DEX pool during the same auction cycle. In 52 cases, one of the tokens swapped is WETH, allowing us to apply the same methodology introduced in the previous section — namely, deriving an implied risk-adjusted CEX price for each bot. To enable direct comparison with observed market data, we further restrict the sample to tokens traded on Binance. This yields 29 auction cycle/DEX pool combinations for which we can compare the bots’ implied prices to contemporaneous Binance prices. We plot the implied DEX price alongside the Binance price for all 29 cases in Appendix 0.A (note that Binance reports prices at one-second intervals, we use the opening price). As expected, the implied DEX prices generally track the Binance price closely. This correlation is particularly strong in cases involving the USDT/WETH and USDC/WETH pairs, which account for the majority of the sample. A weaker but still visible relationship is observed for the UNI/WETH and AAVE/WETH pairs. In contrast, for the DAI/WETH pair, the implied DEX and Binance prices do not exhibit a consistent relationship, likely because Binance is not the primary centralized exchange for DAI trading. This is confirmed by the fact that the total volume on the DAI/ETH market on Binance during our study period was only 2.5 ETH. [^23] For comparison, total volume on Binance during the study period for the other markets was: 1,487 ETH for the USDC/ETH market, 16,209 ETH for the USDT/ETH market, 1,374 ETH for the WBTC/ETH market, 6.9 ETH for the UNI/ETH market, 36 ETH for the AAVE/ETH Binance market.
@@
We complement the visual analysis with a regression approach. For each second $s$, we compute the average implied CEX price and trade volume for Rsync-bot and Titan-bot by considering all trades executed by each bot between seconds s0.5s-0.5 and $s+0.5$. We denote the average implied CEX price by $p_{\text{implied}}$, and associate each observation with timestamp $s$. For technical reasons, we restrict our analysis to the USDT/WETH and USDC/WETH trading pairs. [^24] As discussed earlier, Binance prices for some token pairs appear unreliable. More importantly, the number of observations, values of $p_{\text{implied}}$, and trade volumes vary considerably across tokens. We have 104 observations for WETH/USDC and WETH/USDT combined, but only 24 for WETH/DAI, 18 for WETH/UNI, and 5 or fewer for WBTC/WETH and AAVE/WETH. Moreover, the average value of $p_{\text{implied}}$ is around 3,500 for stablecoin pairs (USDC, USDT, DAI), but drops to 261 for UNI/WETH, 15 for AAVE/WETH, and 0.037 for WBTC/WETH. Trade volumes also vary: average volume is 320.3 ETH for WETH/WBTC, compared to 43.1 ETH for other assets. As a result, separate regressions on the less common pairs are underpowered, while pooled regressions suffer from multicollinearity due to token-specific heterogeneity. This filtering results in 89 total observations: 43 for Rsync-bot and 46 for Titan-bot. We regress $p_{\text{implied}}$ on the contemporaneous Binance opening price, as well as its one- and two-second lag and one-second lead. The regression results, shown in Table 3, indicate that only the contemporaneous and one-second lagged Binance prices are significantly correlated with the bots’ implied CEX prices. Among these, the contemporaneous price has the strongest explanatory power. Therefore, the contemporaneous Binance price serves as the most appropriate benchmark for evaluating $p_{\text{implied}}$.
@@
We then proceed to compare each observation of $p_{\text{implied}}$ to the corresponding contemporaneous Binance price, $p_{\text{binance}}$. To interpret the execution quality from the bot’s perspective, we distinguish between two cases. If a bot sells ETH on-chain, it must be buying ETH on the CEX; in this case, a lower CEX price is better. We therefore define
$p_{\text{diff}}\equiv p_{\text{binance}}-p_{\text{implied}}.$
@@
$p_{\text{diff}}\equiv p_{\text{implied}}-p_{\text{binance}}.$
In both cases, a higher $p_{\text{diff}}$ indicates better execution for the bot. We then calculate
$p_{\text{Improvement}}\equiv\frac{p_{\text{diff}}}{p_{\text{binance}}}\times 100$
as the percentage price improvement relative to the contemporaneous Binance price. The minimum, maximum, and average values of $p_{\text{Improvement}}$ are -0.01428, 0.04655 and 0.01285 respectively.
Table 4 reports the coefficients of two separate regressions. Both include a constant, a dummy variable equal to 1 if the observation is from Titan bot, and time (seconds) since the last block. The difference between the two regressions is that the second includes swap volume (in WETH) in piecewise linear form, where each piece corresponds to 50 ETH. The coefficient on the constant represents the average percentage price improvement ($p_{\text{Improvement}}$) for a zero-volume swap by Rsync-bot, assuming the block was built at time zero in the bidding cycle. The coefficient on Is Titan bot captures the difference in execution quality between Titan-bot and Rsync-bot. The results indicate that the bots achieve better execution than the Binance price, and that Rsync-bot outperforms Titan-bot. Execution quality seems to decline with trade size, although this result needs to be taken with a grain of salt because there are only 2 trades of volume larger than 150. Using the estimated coefficients, we perform back-of-the-envelope calculations. For a swap of volume lower than 100 ETH the implied CEX price for ETH is approximately 2.8 basis points better than the contemporaneous Binance price for both bots.
Table 4: Regression of percentage price improvement over Binance ($p_{\text{Improvement}}$) with and without piecewise volume
@@ 5 Conclusions
We study the Ethereum blockchain by analyzing proposed blocks that were ultimately discarded. To the best of our knowledge, together with the companion paper Pahari and Canidio (2025), ours is the first systematic analysis of non-winning Ethereum blocks. Our findings illuminate key aspects of Ethereum’s block-building process, with implications for ongoing discussions among practitioners and for the academic literature.
While informative, our results should be interpreted in light of certain limitations. As noted, the dataset underrepresents some major builders (notably Beaverbuild) and spans only eight minutes. Moreover, our analysis is primarily descriptive and does not identify the economic mechanisms driving the observed patterns. To address these gaps, we hope that relays, builders, and other participants in the transaction supply chain will make more comprehensive data available in the future, enabling a deeper and more rigorous understanding of how Ethereum is made.
Appendix 0.A Competition between Rsync-bot and Titan-bot on the same pool during the same slot

### v3

@@ 4.4 Competition between Searchers
In the previous section, we highlighted the importance of competition between Rsync-bot and Titan-bot. It turns out that these two searchers compete intensely throughout our entire study period, and their competition is a key determinant of which builder wins the block: every auction cycle for which we have data includes at least one transaction from these bots, which together account for approximately 60% of the builders’ revenues in the blocks where they appear, underscoring their central role in shaping auction outcomes.
@@
We identify 57 instances in which Titan-bot and Rsync-bot compete for execution in the same DEX pool during the same auction cycle. In 52 cases, one of the tokens swapped is WETH, allowing us to apply the same methodology introduced in the previous section — namely, deriving an implied risk-adjusted CEX price for each bot and different moments during a bidding cycle. To enable direct comparison with observed market data, we further restrict the sample to tokens traded on Binance. This yields 29 auction cycle/DEX pool combinations for which we can compare the bots’ implied prices to contemporaneous Binance prices.
We plot the implied DEX price alongside the Binance price for all 29 cases in Appendix A (note that Binance reports prices at one-second intervals, we use the opening price). As expected, the implied DEX prices generally track the Binance price closely. This correlation is particularly strong in cases involving the USDT/WETH and USDC/WETH pairs, which account for the majority of the sample. A weaker but still visible relationship is observed for the UNI/WETH and AAVE/WETH pairs. In contrast, for the DAI/WETH pair, the implied DEX and Binance prices do not exhibit a consistent relationship, likely because Binance is not the primary centralized exchange for DAI trading. This is confirmed by the fact that the total volume on the DAI/ETH market on Binance during our study period was only 2.5 ETH. [^24] For comparison, total volume on Binance during the study period for the other markets was: 1,487 ETH for the USDC/ETH market, 16,209 ETH for the USDT/ETH market, 1,374 ETH for the WBTC/ETH market, 6.9 ETH for the UNI/ETH market, 36 ETH for the AAVE/ETH Binance market.
@@
We complement the visual analysis with a regression approach. For each second $s$, we compute the average implied CEX price and trade volume for Rsync-bot and Titan-bot by considering all trades executed by each bot between seconds s0.5s-0.5 and $s+0.5$. We denote the average implied CEX price for second $s$ by $p_{\text{implied},s}$. For technical reasons, we restrict our analysis to the USDT/WETH and USDC/WETH trading pairs. [^25] As discussed earlier, Binance prices for some token pairs appear unreliable. More importantly, the number of observations, values of $p_{\text{implied},s}$, and trade volumes vary considerably across tokens. We have 104 observations for WETH/USDC and WETH/USDT combined, but only 24 for WETH/DAI, 18 for WETH/UNI, and 5 or fewer for WBTC/WETH and AAVE/WETH. Moreover, the average value of $p_{\text{implied},s}$ is around 3,500 for stablecoin pairs (USDC, USDT, DAI), but drops to 261 for UNI/WETH, 15 for AAVE/WETH, and 0.037 for WBTC/WETH. Trade volumes also vary: average volume is 320.3 ETH for WETH/WBTC, compared to 43.1 ETH for other assets. As a result, separate regressions on the less common pairs are underpowered, while pooled regressions suffer from multicollinearity due to token-specific heterogeneity. This filtering results in 89 total observations: 43 for Rsync-bot and 46 for Titan-bot. We regress $p_{\text{implied},s}$ on the contemporaneous Binance opening price, as well as its one- and two-second lag and one-second lead. The regression results, shown in Table 3, indicate that only the contemporaneous and one-second lagged Binance prices are significantly correlated with the bots’ implied CEX prices. Among these, the contemporaneous price has the strongest explanatory power. Therefore, the contemporaneous Binance price serves as the most appropriate benchmark for evaluating $p_{\text{implied},s}$.
@@
We then proceed to compare each observation of $p_{\text{implied},s}$ to the corresponding contemporaneous Binance price, $p_{\text{binance},s}$. To interpret the execution quality from the bot’s perspective, we distinguish between two cases. If a bot sells ETH on-chain, it must be buying ETH on the CEX; in this case, a lower CEX price is better. We therefore define
$p_{\text{diff},s}\equiv p_{\text{binance},s}-p_{\text{implied},s}.$
@@
$p_{\text{diff},s}\equiv p_{\text{implied},s}-p_{\text{binance},s}.$
In both cases, a higher $p_{\text{diff},s}$ indicates better execution for the bot. We then calculate
$p_{\text{Improvement},s}\equiv\frac{p_{\text{diff},s}}{p_{\text{binance},s}}\times 100$
as the percentage price improvement relative to the contemporaneous Binance price. The minimum, maximum, and average values of $p_{\text{Improvement},s}$ are -0.01428, 0.04655 and 0.01285 respectively.
Table 4 reports the coefficients of two separate regressions. Both include a constant, a dummy variable equal to 1 if the observation is from Titan bot, and time (seconds) since the last block. The difference between the two regressions is that the second includes swap volume (in WETH) in piecewise linear form, where each piece corresponds to 50 ETH. The coefficient on the constant represents the average percentage price improvement ($p_{\text{Improvement},s}$) for a zero-volume swap by Rsync-bot, assuming the block was built at time zero in the bidding cycle. The coefficient on Is Titan bot captures the difference in execution quality between Titan-bot and Rsync-bot. The results indicate that the bots achieve better execution than the Binance price, and that Rsync-bot outperforms Titan-bot. Execution quality seems to improve with trade size, although this result needs to be taken with a grain of salt because there are only 2 trades of volume larger than 150. Using the estimated coefficients, we perform back-of-the-envelope calculations. For a swap of volume lower than 100 ETH the implied CEX price for ETH is approximately 2.8 basis points better than the contemporaneous Binance price for both bots.
Finally, recall that our derivations assume the fee paid by the bots for inclusion equals their risk-adjusted profits expected if their transactions are included on-chain. If this assumption is relaxed, the observed fee provides a lower bound on expected profits: bots may offer builders less than their total expected profits in order to retain part of those profits. In that case, our regression estimates should be interpreted as lower bounds, indicating that the two bots trade on the CEX leg of their arbitrage at prices at least 2.8 basis points better than the contemporaneous Binance price.
Table 4: Regression of percentage price improvement over Binance ($p_{\text{Improvement},s}$) with and without piecewise volume
@@ 5 Conclusions
We study the Ethereum blockchain by analyzing blocks that were proposed for inclusion into the canonical chain but were ultimately discarded. To the best of our knowledge, together with the companion paper pahari2025exclusiveethereumtransactionsevidence, ours is the first systematic analysis of non-winning Ethereum blocks. Our findings illuminate key aspects of Ethereum’s block-building process, with implications for ongoing discussions among practitioners and for the academic literature.
While informative, our results should be interpreted in light of certain limitations. As noted, the dataset underrepresents some major builders (notably Beaverbuild) and spans only eight minutes. To address these gaps, we hope that relays, builders, and other participants in the transaction supply chain will make more comprehensive data available in the future, enabling a deeper and more rigorous understanding of how Ethereum is made.
Appendix A Competition between Rsync-bot and Titan-bot